### v1

@@ Abstract
As modern power systems continue to evolve, accurate power load forecasting remains a critical issue. The phase space reconstruction method can effectively retain the chaotic characteristics of power load from a system dynamics perspective and thus is a promising knowledge-based preprocessing method for power load forecasting. However, limited by its fundamental theory, there is still a gap in implementing a multi-step forecasting scheme in current studies. To bridge this gap, this study proposes a novel multi-step forecasting approach by integrating the PSR with neural networks. Firstly, the useful features in the phase trajectory obtained from the preprocessing of PSR are discussed in detail. Through mathematical derivation, the equivalent characterization of the PSR and another time series preprocessing method, patch segmentation, is demonstrated for the first time. Based on this prior knowledge, an image-based modeling perspective with the global and local feature extraction strategy is introduced. Subsequently, a novel deep learning model, namely PSR-GALIEN, is designed for end-to-end processing, in which the Transformer Encoder and 2D-convolutional neural networks are employed for the extraction of the global and local patterns in the image, and a multi-layer perception based predictor is used for the efficient correlation modeling. Then, extensive experiments are conducted on five real-world benchmark datasets to verify the effectiveness as well as to have an insight into the detailed properties. The results show that, comparing it with six state-of-the-art deep learning models, the forecasting performance of PSR-GALIEN consistently surpasses these baselines, which achieves superior accuracy in both intra-day and day-ahead forecasting scenarios. At the same time, a visualization-based method is proposed to explain the attributions of the forecasting results.
@@ 1 Introduction
An image-based modeling method for the reconstructed phase trajectory is proposed for the first time. This method employs a global and local feature extraction strategy, commonly used in computer vision (CV) tasks, to simultaneously capture corresponding patterns in the 2D image. These patterns represent the overall dynamic evolution characteristics of the trajectory and the temporal locality characteristics within its projection subseries, respectively.
@@
Based on this modeling method, a novel deep-learning model, namely the phase space reconstruction based global and local information enhanced neural network (PSR-GALIEN), is proposed, in which the Transformer Encoder and 2D-convolutional neural network (CNN) are implemented to extract the global and local features in the phase trajectory image respectively, and a well-designed multi-layer perception (MLP) is used as the predictor.
@@
Extensive experiments are conducted in five real-world power load datasets to have an insight to PSR-GALIEN’s performance, and six representative state-of-the-art deep learning models emerged in recent years are deployed for detailed comparison, the results show that PSR-GALIEN substantially exceeds their performance.
@@
The heatmap and regression activation map (RAM) visualization techniques are introduced to explain the extracted global and local features during the inference process, which increases the interpretability of the PSR-GALIEN model.
@@
2 Proposed image-based forecasting approach
@@ 2.1.2 Parameter selection
2.2 Innovative modeling idea
2.2.1 Learning-based forecasting paradigm of PSR
@@ 2.2.2 Data structure and feature analysis
It should be noted that the trade-offs need to be made to consider the way in which this shared data structure is obtained, as the parameters used in the process have practical significance. For example, if the data structure is obtained through PS, and the selected parameters of the segmented stride and the number of segmented patches differ too much from the optimal embedding delay time and the embedding dimension calculated through Eq. 3 and 4, then the dynamic features of the equivalent reconstructed phase trajectories may be diminished, for this parameter selection implemented under the perspective of the PS method does not guarantee the effectiveness of optimal representation of the attractor under the perspective of the PSR method. Thus, decision needs to be given to the prominent characteristics of the time series, for power loads, we mainly consider how to use the method of PSR to extract the chaotic features in it, and at the same time, we can use the prior knowledge of the PS method to extract the temporal locality features from the project sequences in order to bring additional information gain for the forecasting, which we believe is a knowledge gap in the previous studies.
2.2.3 Image-based modeling perspective with global and local feature extraction strategy
In the previous subsections, the data structure and inner features of the phase trajectory is discussed in detail. Thus, considering the coupling and spatial relations of the two different features under the shared matrix data structure, in order to be more flexible in extracting these features from the phase trajectory, and drawing on the feature engineering approaches from the vision Transformer (ViT) [34] and Conformer [39], this paper discards the conventional sequential modelling method for feature extraction and treats it as a two-dimensional grey scale image. At the same time, in this paper, the global and local feature extraction strategy is adopted for the extracting the corresponding features analysed in the previous subsections, which are defined as:
Global Feature: The overall evolution characteristics of phase trajectory in m-dimensional phase space, as shown in the red box in Fig. 1.
Local Feature: The temporal locality characteristics of the projection sequences of the phase trajectory in each dimension (equivalently, the segmented patch subseries), as shown in the green boxes in Fig. 1.
Taking the above analyses together, it is now clear what kind of features can be extracted from the phase trajectories, and in the next section, the way to build a deep learning network to conduct such feature extraction as well as to form the end-to-end correlation model will be introduced.

### v2

@@ Abstract
As modern power systems continue to evolve, accurate power load forecasting remains a critical issue in energy management. The phase space reconstruction (PSR) method can effectively retain the inner chaotic property of power load from a system dynamics perspective and thus is a promising knowledge-based preprocessing method for power load forecasting. In order to fully utilize the capability of PSR method to model the non-stationary characteristics within power load, and to solve the problem of the difficulty in applying traditional PSR prediction methods to form a general multi-step forecasting scheme, this study proposes a novel multi-step forecasting approach by delicately integrating the PSR with neural networks to establish an end-to-end learning system. Firstly, the useful features in the phase trajectory obtained from the preprocessing of PSR are discussed in detail. Through mathematical derivation, the equivalent characterization of the PSR and another time series preprocessing method, patch segmentation (PS), is demonstrated for the first time. Based on this prior knowledge, an image-based modeling perspective with the global and local feature extraction strategy is introduced. Subsequently, a novel deep learning model, namely PSR-GALIEN, is designed for end-to-end processing, in which the Transformer Encoder and 2D-convolutional neural networks (CNN) are employed for the extraction of the global and local patterns in the image, and a multi-layer perception (MLP) based predictor is used for the efficient correlation modeling. Then, extensive experiments are conducted on five real-world benchmark datasets to verify the effectiveness as well as to have an insight into the detailed performance of the PSR-GALIEN. The results show that, compared with six state-of-the-art deep learning models, the forecasting performance of PSR-GALIEN consistently surpasses these baselines, achieving superior accuracy in both intra-day and day-ahead forecasting scenarios. At the same time, the attributions of its forecasting results can be explained through the visualization-based method, which significantly increases the interpretability.
@@ 1 Introduction
A novel image-based modeling and forecasting framework is proposed to make full utilization of the PSR method in feature engineering, which adopts a global and local feature extraction strategy to simultaneously extract the corresponding patterns in the phase trajectory image, and their correlations to the future target series are established jointly.
@@
Based on this framework, a novel deep-learning model, namely PSR-GALIEN, is proposed for efficient end-to-end processing, in which the Transformer Encoder and 2D-convolutional neural network (CNN) are implemented to extract the global and local features in this image respectively, and a multi-layer perception (MLP) based predictor is used for correlation modeling.
@@
Extensive experiments are conducted to have a comprehensive insight to PSR-GALIEN’s performance, demonstrating the effectiveness of this image-based approach, as well as the superior forecasting accuracy of PSR-GALIEN in the real-world intra-day and day-ahead forecasting scenarios.
@@
The feature interpretation method of using the regression activation map (RAM) and heatmap to explain the utilized features of PSR-GALIEN in the forecasting process is proposed, which is critical for understanding the attribution of the predictions results.
@@
2 Proposed image-based methodology
@@ 2.1.2 Parameter selection
2.2 Image-based modeling and forecasting framework
2.2.1 General Learning-based forecasting paradigm of PSR
@@ 2.2.2 Data structure and feature analysis
It should be noted that the trade-offs need to be made to consider the way in which this shared data structure is obtained, as the parameters used in the process have practical significance, and decision needs to be given to the prominent characteristics of the time series. For power loads, we mainly consider how to use the method of PSR to extract the chaotic features in it, and at the same time, we can use the PS method to extract the temporal locality features from the project sequences in order to bring additional information gain for the forecasting, which we believe is a knowledge gap in the previous studies.
2.2.3 Feature extraction and Correlation modeling strategy
In the previous subsections, the data structure and inner features of the phase trajectory is discussed in detail. Thus, considering the coupling and spatial relations of the two different features under the shared matrix data structure, in order to be more flexible in extracting these features from the phase trajectory, and drawing on the feature engineering approaches from the vision Transformer (ViT) [34] and Conformer [39], this paper discards the conventional sequential modelling method for feature extraction and treats it as a two-dimensional greyscale image. At the same time, the global and local feature extraction strategy is adopted for the extracting the corresponding features analysed in the previous subsections, which are defined as:
Global Feature: The overall evolution characteristics of phase trajectory in m-dimensional phase space, which can be characterized by the global pattern of the image, as shown in the red box in Fig. 1.
Local Feature: The temporal locality characteristics of the projection sequences of the phase trajectory in each dimension (equivalently, the segmented patch subseries), which can be characterized by the local patterns of the image, as shown in the green boxes in Fig. 1.
Taking the above analyses together, it is now clear what kind of features can be extracted from the phase trajectories, intuitively, both the two features can be extracted and utilized to establish a correlation model, but the connection between them is unrealised. Thus, based on this prior knowledge which is demonstrated in the subsections, we consider to form a jointly correlation model to make the best of this data structure.
In the next section, the way to build a deep learning architecture to conduct such feature extraction as well as feature correlation strategy to form an end-to-end learning system will be introduced.